If U(R) denotes the group of units of a ring R, show that

U(R1 × R2) = U(R1) × U(R2) for rings R1 and R2.


1
Expert's answer
2021-10-08T04:44:51-0400

Let 1 denote the unity element in each of the rings Ri. So, (1, 1) is the unity element in R1"\times"R2, since if (a1, a2) ∈ R1"\times"R2, then

(a1, a2)(1, 1) = (a1 · 1, a2 · 1) = (a1, a2),

and

(1, 1)(a1, a2) = (1 · a1, 1 · a2) = (a1, a2). 

Now, suppose that (u1, u2) ∈ U(R1"\times"R2). Then, since this element is a unit in the ring R1"\times"R2, then by definition, this means there exists an element (x1, x2) ∈ R1"\times"R2 such that 

(u1, u2)(x1, x2) = (1, 1). 

But we have, by definition of multiplication in the direct sum of rings,

(u1, u2)(x1, x2) = (u1x1, u2x2)

 That is, we have

(u1, u2)(x1, x2) = (u1x1, u2x2) = (1, 1)

So, for each i, 1 ≤ i ≤ n, we have uixi = 1 = xiui . That is, xi ∈ Ri and xi is a multiplicative inverse to ui . Thus ui ∈ U(Ri). Therefore, (u1, u2) ∈ U(R1)"\times"U(R2). Therefore, we have

U(R1"\times"R2) ⊆ U(R1)"\times"U(R2).


To prove the other direction of containment, suppose that (y1, y2) ∈ U(R1)"\times"U(R2). This means

yi ∈ U(Ri). So, there exists a wi ∈ Ri such that yiwi = 1 = wiyi .

Consider the element (w1, w2) ∈ R1"\times"R2. Then we have

(y1, y2)(w1, w2) = (y1w1, y2w2) = (1, 1),

and

(w1, w2)(y1, y2) = (w1y1, w2y2) = (1, 1).

Thus, (w1, w2) is a multiplicative inverse for the element (y1, y2) in R1"\times"R2. In particular, we have (y1, y2) ∈ U(R1)"\times"U(R2). Therefore,

U(R1)"\times"U(R2) ⊆ U(R1"\times"R2).

Thus, we have now shown U(R1"\times"R2) = U(R1)"\times"U(R2).